OpenStudy (anonymous):

What is the solution to the open sentence? 4y – 3 = y + 6

OpenStudy (anonymous):

Get your y's to one side \[4y-3=y+6\] subtract y frome the right because we don't like negative numbers

OpenStudy (anonymous):

\[3y-3=6\] now move the minus 3 to the right (still don't like negative numbers)

OpenStudy (anonymous):

\[3y=9\](because you had to add the 3 to cancel it out on the left)

OpenStudy (anonymous):

And now simply divide by 3

OpenStudy (anonymous):

Thanks c:

OpenStudy (anonymous):

You are most welcome